OpenStudy (anonymous):

3 cards are drawn from a standard deck. what are the possibilities of getting 3 kings?

OpenStudy (anonymous):

You mean what is the probability? 4/52 x 3/51 x 2/50 = 1 / 5525 or 0.000181

OpenStudy (anonymous):

12 / 52 all cards = 52 and kings = 12 so, 12/52

OpenStudy (anonymous):

Not right, @rukhsar . Pr(1st card is a king) * Pr(2nd card is a king) * Pr(3rd card is a king) After the first card is drawn, there are 51 cards left, 3 of which are kings. After the second card is drawn, there are 50 cards left, 2 of which are kings.

OpenStudy (anonymous):

yes. thanks BangkokGarrett

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

@BangkokGarrett but its about 3 cards,, so 3x4 = 12 am i right??

OpenStudy (anonymous):

There are 3 events happening that you have to calculate the probabilities of separately. Pr(all 3 are kings) = Pr(1st card is King) * Pr(2nd card is a king) * Pr(3rd card is a king) 4/52 * 3/51 * 2/50

OpenStudy (anonymous):

There aren't 12 cards in the deck that are kings. There are only 4 kings in the deck.

OpenStudy (anonymous):

12/52 would be the probability of drawing one card and getting a Jack or a Queen or a King

OpenStudy (anonymous):

okayyy... thanks!